A boat can travel on a river 7 miles upstream in the same time it takes to travel 9 miles downstream. Find the speed of the current if the speed of the boat in still water is 8 miles per hour.
Answer by TimothyLamb(4379)

---
x = boat speed
y = current speed
x = 8 mph
---
s = d/t
t = d/s
---
t = 7/(x - y)
t = 9/(x + y)
---
7/(x - y) = 9/(x + y)
7(x + y) = 9(x - y)
7x + 7y = 9x - 9y
9y + 7y = 9x - 7x
16y = 2x
y = 2x/16
y = (2*8)/16
---
y = current speed = 1 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php